# Binarized Simplicial Convolutional Neural Networks

### v1

@@ 4.2 Over-Smoothing Alleviation via Architectural Simplifications
4.3 Computational Complexity Reduction via Binarization of Simplicial Features
Let us now examine how to enhance the time efficiency of simplicial convolution while maintaining the effectiveness of its simplicial embedding. One of the many solutions to this is the Binarization similar to what is seen in the Bi-GCN [27], where both the trainable weights and the features are binarized and the forward propagation is done by binary operations bit count and XOR to combine the binarized weights and features into a new node embedding. Another approach is seen in the Graph-Sign algorithms [28], which delivers a sign-error update at each iteration that has a low run time and a low complexity. Aside from the fact that these node-based algorithms are not suitable to be deployed on simplicial complexes, there are a few additional problems to be addressed. The Bi-GCN binarizes both the weights and features, leading to reduced performance under certain circumstances. Also, even though Bi-GCN uses binary operations to optimize the computation cost, the forward propagation is indifferent and mathematically identical compared with the original GCN. The Graph-Sign algorithm operates on a predefined bandlimited filter, which may require prior expert knowledge to obtain.
We propose a new forward propagation strategy where we combine feature normalization, feature binarization, and binary-sign propagation together to form a fast and accurate Binarized Simplicial Convolution. Let us first begin with the multi-variate simplicial convolution in (10). Following the graph-based feature binarization seen in Bi-GCN [27], we define the feature binarization of simplicial data $\mathbf{X}_{k}$ on $k$-simplex $\mathcal{S}_{k}$ as
where $\circ$ is the Hadamard product, Sign()()( ) is the Sign function and $\mathbf{\bar{X}_{k}}$ is the (weighted) binarized simplicial features. The matrix $\mathbf{M}_{k}$ acts as the weights for the binarized features and is formed by a feature normalization:
${m_{i}}=\frac{norm_{1}(\mathbf{r}_{k})}{d_{k}},$ (18)
where $norm_{1}(\mathbf{r}_{k})$ is the $l_{1}$-norm done on the $i^{th}$ row of $\mathbf{X}_{k}$ ($\mathbf{r}_{k}$ in (18)) and $d_{k}$ is the number of features in $\mathcal{S}_{k}$. All elements of the same row in $\mathbf{M}_{k}$ have the same value, but this is only for mathematical representation convenience. In practice, $\mathbf{M}_{k}$ is stored as a vector $\boldsymbol{m}_{k}$ and is broadcasted using element-wise multiplications when used to multiply with $\text{Sign}(\mathbf{X}_{k})$. Notice that the Sign() function only outputs binary values $-1$ and $+1$ assuming that we ignore the zeros, this means that the size of the binarized features $\mathbf{\bar{X}_{k}}$ is reduced in practice when compared to $\mathbf{X}_{k}$ [27].
$\mathbf{\bar{Z}}_{k,{p+1}}=\sigma\left(\mathbf{L}_{k,l}\mathbf{\bar{Z}}_{k,{p}}\mathbf{\Gamma}_{k}+\mathbf{L}_{k,u}\mathbf{\bar{Z}}_{k,{p}}\mathbf{\Theta}_{k}+\mathbf{\bar{Z}}_{k,{p}}\mathbf{\Xi}_{k}\right).$ (19)
Looking at the expression in (19), aside from benefiting from the model size reduction, it still follows the exact same forward propagation strategy as SCNN, making it have roughly the same time to execute in practice. We would like to further improve the run speed of this Binary setup by propagating only the Sign() part. Using the two properties of Hadamard products $\mathbf{A}\circ\mathbf{B}=\mathbf{B}\circ\mathbf{A}$ and $\mathbf{A}\circ\mathbf{(B+C)}=\mathbf{A}\circ\mathbf{B}+\mathbf{A}\circ\mathbf{C}$, with some algebraic derivations, equation (19) becomes
$\begin{split}\mathbf{\bar{Z}}_{k,{p+1}}=&\mathbf{M}_{k,p}\circ(\mathbf{L}_{k,l}\text{Sign}(\mathbf{\bar{Z}}_{k,{p}})\mathbf{\Gamma}_{k}+\\ &\mathbf{L}_{k,u}\text{Sign}(\mathbf{\bar{Z}}_{k,{p}})\mathbf{\Theta}_{k}+\text{Sign}(\mathbf{\bar{Z}}_{k,{p}})\mathbf{\Xi}_{k}).\end{split}$ (20)
Notice that the $l_{1}$-norm is nonnegative, so at each layer, $\mathbf{M}_{k,p}$ can be viewed as $\mathbf{M}_{k,p}=$ ReLu($\mathbf{M}_{k,p}$). The input of the $p+1^{th}$ layer in (20) will binarize the output of the $p^{th}$ layer regardless of the magnitude, so we can safely input only the binarized simplicial features (without the weights $\mathbf{M}_{k}$) to the $p+1^{th}$ layer when $p>1$. Also, we should point out that the Sign() function is non-linear. In this way, the nonlinearity in both the Sign() function and the $l_{1}$-norm can serve as activations in forward propagation. The forward propagation of a series of $P$ concatenated Bi-SCNN is obtained by rearranging (20) a using the property $\mathbf{A}\circ\mathbf{B}=\mathbf{B}\circ\mathbf{A}$ again:
$\mathbf{\bar{Z}}_{k,{P}}=\mathbf{M}_{k,p}\circ\mathbf{M}_{k,P-1}\circ...\circ\mathbf{M}_{k,1}\circ\mathbf{Q}_{k,P-1},$ (21)
@@ 4.3 Computational Complexity Reduction via Binarization of Simplicial Features
$\mathbf{Q}_{k,p}=\text{Sign}(\mathbf{L}_{k,l}\mathbf{\bar{Z}}_{k,p-1}\mathbf{\Gamma}_{k}+\mathbf{L}_{k,u}\mathbf{\bar{Z}}_{k,p-1}\mathbf{\Theta}_{k}+\mathbf{\bar{Z}}_{k,p-1}\mathbf{\Xi}_{k}).$ (22)
@@
Even though the Sign() function has no derivative at 0 due to its discontinuity at 00 and a derivative of 00 for all other values, we approximate Sign() using a hard tanh function, allowing us to conduct forward propagation and backward propagation as usual. The way to achieve the forward propagation in (21) in practice is that each Bi-SCNN layer will have two outputs: the feature normalization $\mathbf{M}_{k,p}$ and feature binarization $\mathbf{Q}_{k,p}$. Notice that only $\mathbf{Q}_{k,p}$ will be forward-propagated to the next Bi-SCNN layer, meaning that $\mathbf{Q}_{k,p}=\mathbf{Z}_{k,p}$. Exceptions are: at the first layer where $\mathbf{Z}_{k,0}=\mathbf{X}_{k}$ is the input, and the final layer all the $\mathbf{M}_{k,p}$ are multiplied together at the end to form $\mathbf{\bar{Z}}_{k,{P}}$. An illustration of a Bi-SCNN network with $P=3$ is shown in Fig. 2.
@@
Intuitively, a single Bi-SCNN layer on a $k$-simplex can be understood as using the (upper and lower) Hodge Laplacians to aggregate the weighted simplicial features with the addition of the harmonic component. From the spatial Topological Signal Processing perspective, the weights are trained to optimize a binary-sign feature aggregation with a matrix that records the magnitude of each aggregation. From the spectral Topological Signal Processing perspective, the goal of training a Bi-SCNN is to obtain a filter for the binary-sign update that best fits the spectrum of the training data. An illustration of one Bi-SCNN layer can be found in Fig. 2.
@@
Aside from the fact that all algorithms benefit from the sparsity of $\mathbf{L}_{k}$, let us look into (21) to see the high time-efficiency of Bi-SCNN. In practice, Sign() is approximated by Hardtanh(), meaning that a percentage $r_{k}$ of the matrix $\mathbf{\bar{Z}}_{k,{p}}$ consists only of $1$s and $-1$s. In this case, $r_{k}$ element-wise multiplications can be omitted in the matrix multiplication, reducing $N_{k}^{2}d_{k}$ multiplications to $(1-r_{k})N_{k}^{2}d_{k}$ multiplications. Assuming each multiplication is 1 FLOP, each Bi-SCNN layer is at least $r_{k}$ FLOPS faster than the SCNN layer for the $2^{nd}$ layer and beyond. Also, in the actual implementation of SCNN, to compute $\mathbf{Z}_{k}^{p}$, $\mathbf{L}_{k,l}$, $\mathbf{L}_{k,u}$, and $\mathbf{I}_{k}$ are concatenated into a $3N_{k}\times N_{k}$ matrix for each $j=0...J$; then $J$ matrices are stacked into a $d_{p}\times N_{k}\times J$ tensor. Trainable weights $\mathbf{\Theta}_{k}$, $\Gamma_{k,j}$, and $\Xi_{k}$ are stacked into a $d_{k,p}\times d_{k,p+1}\times J$ tensor in a similar manner. Afterward, $\mathbf{Z}_{k,p+1}$ is computed by the Einstein summation between two tensors (aggregation and weight tensors) of size $d_{p}\times N_{k}\times J$ and $d_{k,p}\times d_{k,p+1}\times J$ back into a matrix of size $N_{k}\times d_{p+1}$. This matrix concatenation, stacking, and Einstein summation approach of the SCNN is less time efficient than direct matrix multiplication in Bi-SCNN considering we are using binarized features, especially when $J>1$. We should emphasize again that $\mathbf{M}_{k,p}$ is a broadcasted vector in practice, making the Bi-SCNN more efficient.
@@ 5 Experiment Results and Discussion
The Citation Complex [15] and the Ocean Drifter Complex [14] are chosen because both datasets are already preprocessed to have features defined in simplexes with order $k\geq 1$, making them suitable choices to test the performance of an algorithm on simplicial complexes. We would like to verify that the Bi-SCNN has a faster execution time while maintaining good accuracy when compared against existing simplicial algorithms. The experiments are conducted on one Nvidia RTX 3090 GPU with PyTorch version 1.8.0 and CUDA version 11.1. None of the two datasets provided a validation set so the experiments are tuned based on the training performance. All the experiments were repeated 10 times; we calculated the mean and standard derivation of the metrics. A visualization of the ocean drifter complex with illustrations of 5 different ocean drifter trajectories is shown in Figure 4.
SNN-2 SCNN-2 SCNN-3 Bi-SCNN-3 Bi-SCNN-2
@@
$\mathcal{S}_{0}$ 90.60$\pm$03.47 90.51$\pm$0.34 90.57$\pm$0.38 84.52$\pm$17.95 90.65$\pm$0.32
$\mathcal{S}_{1}$ 91.00$\pm$02.45 91.02$\pm$0.25 90.52$\pm$1.43 90.77$\pm$0.24 91.03$\pm$0.24
$\mathcal{S}_{2}$ 90.94$\pm$05.59 90.60$\pm$0.67 91.22$\pm$0.16 87.81$\pm$01.91 91.22$\pm$0.16
$\mathcal{S}_{3}$ 91.45$\pm$03828 91.37$\pm$0.14 91.58$\pm$0.12 89.74$\pm$01.65 91.58$\pm$0.12
$\mathcal{S}_{4}$ 87.89$\pm$11.59 91.65$\pm$0.56 91.88$\pm$0.23 85.04$\pm$16.80 91.92$\pm$0.20
$\mathcal{S}_{5}$ 83.61$\pm$18.65 92.07$\pm$0.33 82.96$\pm$27.64 89.44$\pm$0.49 92.21$\pm$0.23
@@
$\mathcal{S}_{0}$ 81.22$\pm$0.63 81.36$\pm$0.64 81.14$\pm$0.63 81.08$\pm$0.61 81.39$\pm$0.66
$\mathcal{S}_{1}$ 82.17$\pm$0.32 82.16$\pm$0.32 69.76$\pm$26.24 82.07$\pm$0.25 82.20$\pm$0.31
$\mathcal{S}_{2}$ 82.42$\pm$0.49 82.26$\pm$0.51 74.42$\pm$24.81 80.53$\pm$0.58 82.68$\pm$0.27
$\mathcal{S}_{3}$ 83.11$\pm$0.38 82.97$\pm$0.26 83.23$\pm$0.15 81.81$\pm$1.28 83.23$\pm$0.15
$\mathcal{S}_{4}$ 75.63$\pm$23.65 83.47$\pm$0.40 81.84$\pm$5.39 82.79$\pm$0.45 83.64$\pm$0.19
$\mathcal{S}_{5}$ 73.03$\pm$25.38 84.22$\pm$0.27 75.88$\pm$25.30 66.45$\pm$2.69 84.34$\pm$0.23
@@
$\mathcal{S}_{0}$ 72.16$\pm$0.66 72.27$\pm$0.57 71.82$\pm$1.28 72.05$\pm$0.61 72.33$\pm$0.50
$\mathcal{S}_{1}$ 73.15$\pm$0.42 73.13$\pm$0.44 72.87$\pm$0.52 72.64$\pm$0.79 73.98$\pm$0.41
$\mathcal{S}_{2}$ 73.73$\pm$0.67 73.46$\pm$0.51 66.70$\pm$21.96 72.20$\pm$0.32 73.98$\pm$0.26
$\mathcal{S}_{3}$ 74.70$\pm$0.30 74.52$\pm$0.21 67.29$\pm$22.43 72.44$\pm$0.40 74.78$\pm$0.16
$\mathcal{S}_{4}$ 68.31$\pm$21.98 75.56$\pm$0.26 68.05$\pm$22.69 74.73$\pm$08.27 75.68$\pm$0.20
$\mathcal{S}_{5}$ 72.09$\pm$10.87 76.52$\pm$0.31 76.63$\pm$0.24 65.00$\pm$18.89 76.63$\pm$0.23
@@
$\mathcal{S}_{0}$ 62.36$\pm$0.72 62.78$\pm$0.65 62.07$\pm$1.83 62.59$\pm$0.61 62.81$\pm$0.65
$\mathcal{S}_{1}$ 64.10$\pm$0.28 64.10$\pm$0.28 63.74$\pm$0.91 63.57$\pm$0.79 64.10$\pm$0.28
$\mathcal{S}_{2}$ 64.73$\pm$0.79 64.54$\pm$0.63 65.52$\pm$0.20 63.86$\pm$0.32 65.17$\pm$0.20
$\mathcal{S}_{3}$ 66.26$\pm$0.28 66.07$\pm$0.33 66.31$\pm$0.21 65.40$\pm$0.40 66.31$\pm$0.21
$\mathcal{S}_{4}$ 66.89$\pm$1.66 67.33$\pm$0.39 50.26$\pm$27.14 66.85$\pm$0.83 67.57$\pm$0.23
$\mathcal{S}_{5}$ 61.98$\pm$14.03 68.86$\pm$0.18 62.21$\pm$20.31 53.91$\pm$18.89 68.98$\pm$0.18
@@
$\mathcal{S}_{0}$ 54.01$\pm$0.5457 54.03$\pm$0.06 53.98$\pm$0.58 50.28$\pm$5.35 54.18$\pm$0.52
$\mathcal{S}_{1}$ 55.05$\pm$0.3947 55.08$\pm$0.04 54.33$\pm$1.37 54.21$\pm$0.94 55.81$\pm$0.40
$\mathcal{S}_{2}$ 55.32$\pm$2.1300 55.78$\pm$0.06 56.27$\pm$0.24 52.95$\pm$4.04 56.27$\pm$0.24
$\mathcal{S}_{3}$ 57.81$\pm$0.3062 57.33$\pm$0.03 57.82$\pm$0.28 57.19$\pm$0.37 57.85$\pm$0.25
$\mathcal{S}_{4}$ 57.64$\pm$4.4258 59.05$\pm$0.04 59.54$\pm$0.29 57.98$\pm$3.21 59.54$\pm$0.29
$\mathcal{S}_{5}$ 49.87$\pm$16.5068 60.72$\pm$0.09 61.15$\pm$0.19 55.39$\pm$6.39 61.15$\pm$0.19
@@
SNN-2 SCNN-2 SCNN-3 Bi-SCNN-3 Bi-SCNN-2

### v2

@@ 4.4 Simplicial Feature Binarization
Figure 6: An illustration of a $P$-layer Bi-SCNN network in (26) for simplicial order $K=2$ and $P=3$. The forward propagation is indicated in the dashed arrows with each colored path corresponding to the component of the same color in Figure 5. For the input $\mathbf{X}$, the adjacencies within a simplicial signal $\mathbf{X}_{k}$ are denoted in black.
The forward propagation of a series of $P$ concatenated Bi-SCNN is obtained by rearranging (25) a using the property $\mathbf{A}\circ\mathbf{B}=\mathbf{B}\circ\mathbf{A}$ again:
$\mathbf{\bar{Z}}_{k,{P}}=\mathbf{M}_{k,P}\circ\mathbf{M}_{k,P-1}\circ...\circ\mathbf{M}_{k,2}\circ\mathbf{Q}_{k,P-1},$ (26)
@@
$\mathbf{Q}_{k,p}=\text{Sign}(\mathbf{L}_{k,l}\mathbf{\bar{Z}}_{k,p}\mathbf{\Gamma}_{k,{p}}+\mathbf{L}_{k,u}\mathbf{\bar{Z}}_{k,p}\mathbf{\Theta}_{k}+\mathbf{\bar{Z}}_{k,p}\mathbf{\Xi}_{k}).$ (27)
@@
The way to achieve the forward propagation of Bi-SCNN in (26) in practice is that each Bi-SCNN layer will have two outputs: the feature normalization $\mathbf{M}_{k,p}$ and feature binarization $\mathbf{Q}_{k,p}$. Within the two outputs, notice that only $\mathbf{Q}_{k,p}$ will be forward-propagated to the next Bi-SCNN layer, meaning that $\mathbf{Q}_{k,p}=\mathbf{\bar{Z}}_{k,p}$, with the following two exceptions: 1. at the first layer where $\mathbf{X}_{k}$ is the input, and 2. the final layer all the $\mathbf{M}_{k,p}$ are multiplied together at the end to form $\mathbf{\bar{Z}}_{k,{P}}$. Each of the normalization $\mathbf{M}_{k,p}$ is propagated straight to the final Bi-SCNN layer and is combined using element-wise multiplication. An illustration of one Bi-SCNN layer is shown in Fig.5, and an illustration of a Bi-SCNN network with $P=3$ is shown in Fig.6.
4.5 Interpretation and Analysis of Bi-SCNN
@@
Intuitively, a single Bi-SCNN layer on a $k$-simplex can be understood as using the (upper and lower) Hodge Laplacians to aggregate the weighted simplicial features with the addition of the harmonic component. From the spatial TSP perspective, the weights are trained to optimize a binary-sign feature aggregation with a matrix that records the magnitude of each aggregation. From the spectral TSP perspective, the goal of training a Bi-SCNN is to obtain a filter for the binary-sign update that best fits the spectrum of the training data.
@@
Aside from the fact that all simplicial-convolution-based algorithms benefit from the sparsity of $\mathbf{L}_{k}$, let us look into (26) to see the high time-efficiency of Bi-SCNN. In practice, Sign() is approximated by a hard tanh function, meaning that a percentage $r_{k}$ of the matrix $\mathbf{\bar{Z}}_{k,{p}}$ consists only of $1$s and $-1$s. In this case, $r_{k}$ element-wise multiplications can be omitted in the matrix multiplication, reducing $N_{k}^{2}d_{k}$ multiplications to $(1-r_{k})N_{k}^{2}d_{k}$ multiplications. Assuming each multiplication is 1 FLOP, each Bi-SCNN layer is at least $r_{k}$ FLOPS faster than the SCNN layer for the $2^{nd}$ layer and beyond.
Even though the Sign() used in the Bi-SCNN layers (26) has no derivative at 0 due to its discontinuity at 00 and a derivative of 00 for all other values, we approximate Sign() using a hard tanh function, allowing us to conduct forward propagation and backward propagation as usual. This approach is known as the straight-through estimator, where gradient-based optimization is enabled on non-differentiable functions by using a proxy gradient during backward propagation (Yin et al.,, 2019; Le et al.,, 2022).
Let us compare the model complexity of Bi-SCNN with its predecessors. In the actual implementation of SCNN, to compute $\mathbf{Z}_{k}^{p}$, $\mathbf{L}_{k,l}$, $\mathbf{L}_{k,u}$, and $\mathbf{I}_{k}$ are concatenated into a $3N_{k}\times N_{k}$ matrix for each $j=0...J$; then $J$ matrices are stacked into a $d_{p}\times N_{k}\times J$ tensor. Trainable weights $\mathbf{\Theta}_{k}$, $\Gamma_{k,j}$, and $\Xi_{k}$ are stacked into a $d_{k,p}\times d_{k,p+1}\times J$ tensor in a similar manner. Afterward, in SCNN, $\mathbf{Z}_{k,p+1}$ is computed by the Einstein summation between two tensors (aggregation and weight tensors) of size $d_{p}\times N_{k}\times J$ and $d_{k,p}\times d_{k,p+1}\times J$ back into a matrix of size $N_{k}\times d_{p+1}$. This matrix concatenation, stacking, and Einstein summation approach of the SCNN is less time efficient than direct matrix multiplication in Bi-SCNN considering we are using binarized features, especially when $J>1$. The SNN in (12) has a similar model complexity to the SCNN in (13) because instead of splitting the Hodge Laplacian matrix $\mathbf{L}_{k}$ using Hodge decomposition, SNN kept the Hodge Laplacian matrix as a whole. However, it still has a length $J$ component in each SNN layer. Compared to SCNN and Bi-SCNN, the SNN has lower representation power because SNN does not adopt the Hodge decomposition in (7) to model upper, lower, and harmonic components separately. Additionally, SNN has high computation and memory demands similar to the SCNN.
On the other hand, we should emphasize again that $\mathbf{M}_{k,p}$ is implemented using a vector in practice, followed by element-wise multiplication instead of matrix multiplication, making the Bi-SCNN more time-efficient. Lastly, we have mentioned earlier that the Sign() function is being approximated by a hard tanh function, which is a straight-through estimator, allowing the backward propagation calculation to be extremely efficient as straight-through estimators rely only on low-computational complexity or predefined values for the gradient calculation.
@@ 5 Experiment Results and Discussion
The Citation Complex (Ebli et al.,, 2020) and the Ocean Drifter Complex (Schaub et al.,, 2020) are chosen because both datasets are already preprocessed to have features defined in simplexes with order $k\geq 1$, making them suitable choices to test the performance of an algorithm on simplicial complexes. We would like to verify that the Bi-SCNN has a faster execution time while maintaining good accuracy when compared against existing simplicial algorithms. The experiments are conducted on one Nvidia RTX 3090 GPU with PyTorch version 1.8.0 and CUDA version 11.1. None of the two datasets provided a validation set so the experiments are tuned based on the training performance. All the experiments were repeated 10 times; we calculated the mean and standard derivation of the metrics. A visualization of the ocean drifter complex with illustrations of 5 different ocean drifter trajectories is shown in Figure 9.
Figure 7: Mean run time of each algorithm on the Citation Complex for $10\%-50\%$ missing data.
SNN-2 SCNN-2 SCNN-3 Bi-SCNN-3 Bi-SCNN-2 MPNN-2
@@
$\mathcal{S}_{0}$ 90.60$\pm$3.47 90.51$\pm$0.34 90.57$\pm$0.38 84.52$\pm$17.95 90.65$\pm$0.32 6.13$\pm$3.07
$\mathcal{S}_{1}$ 91.00$\pm$02.45 91.02$\pm$0.25 90.52$\pm$1.43 90.77$\pm$0.24 91.03$\pm$0.24 72.84$\pm$36.42
$\mathcal{S}_{2}$ 90.94$\pm$05.59 90.60$\pm$0.67 91.22$\pm$0.16 87.81$\pm$01.91 91.22$\pm$0.16 36.50$\pm$44.70
$\mathcal{S}_{3}$ 91.45$\pm$3.828 91.37$\pm$0.14 91.58$\pm$0.12 89.74$\pm$1.65 91.58$\pm$0.12 38.17 $\pm$ 43.92
$\mathcal{S}_{4}$ 87.89$\pm$11.59 91.65$\pm$0.56 91.88$\pm$0.23 85.04$\pm$16.80 91.92$\pm$0.20 38.16 $\pm$ 40.48
$\mathcal{S}_{5}$ 83.61$\pm$18.65 92.07$\pm$0.33 82.96$\pm$27.64 89.44$\pm$0.49 92.21$\pm$0.23 22.17 $\pm$ 0.00
@@
$\mathcal{S}_{0}$ 81.22$\pm$0.63 81.36$\pm$0.64 81.14$\pm$0.63 81.08$\pm$0.61 81.39$\pm$0.66 3.07 $\pm$ 3.76
$\mathcal{S}_{1}$ 82.17$\pm$0.32 82.16$\pm$0.32 69.76$\pm$26.24 82.07$\pm$0.25 82.20$\pm$0.31 49.38 $\pm$ 40.14
$\mathcal{S}_{2}$ 82.42$\pm$0.49 82.26$\pm$0.51 74.42$\pm$24.81 80.53$\pm$0.58 82.68$\pm$0.27 62.33 $\pm$ 33.15
$\mathcal{S}_{3}$ 83.11$\pm$0.38 82.97$\pm$0.26 83.23$\pm$0.15 81.81$\pm$1.28 83.23$\pm$0.15 33.27 $\pm$ 40.74
$\mathcal{S}_{4}$ 75.63$\pm$23.65 83.47$\pm$0.40 81.84$\pm$5.39 82.79$\pm$0.45 83.64$\pm$0.19 18.33 $\pm$ 32.92
$\mathcal{S}_{5}$ 73.03$\pm$25.38 84.22$\pm$0.27 75.88$\pm$25.30 66.45$\pm$2.69 84.34$\pm$0.23 22.17 $\pm$ 0.00
@@
$\mathcal{S}_{0}$ 72.16$\pm$0.66 72.27$\pm$0.57 71.82$\pm$1.28 72.05$\pm$0.61 72.33$\pm$0.50 6.90 $\pm$ 2.30
$\mathcal{S}_{1}$ 73.15$\pm$0.42 73.13$\pm$0.44 72.87$\pm$0.52 72.64$\pm$0.79 73.98$\pm$0.41 21.91 $\pm$ 33.44
$\mathcal{S}_{2}$ 73.73$\pm$0.67 73.46$\pm$0.51 66.70$\pm$21.96 72.20$\pm$0.32 73.98$\pm$0.26 49.79 $\pm$ 33.08
$\mathcal{S}_{3}$ 74.70$\pm$0.30 74.52$\pm$0.21 67.29$\pm$22.43 72.44$\pm$0.40 74.78$\pm$0.16 37.48 $\pm$ 34.26
$\mathcal{S}_{4}$ 68.31$\pm$21.98 75.56$\pm$0.26 68.05$\pm$22.69 74.73$\pm$08.27 75.68$\pm$0.20 30.47 $\pm$ 36.92
$\mathcal{S}_{5}$ 72.09$\pm$10.87 76.52$\pm$0.31 76.63$\pm$0.24 65.00$\pm$18.89 76.63$\pm$0.23 22.17 $\pm$ 0.00
@@
$\mathcal{S}_{0}$ 62.36$\pm$0.72 62.78$\pm$0.65 62.07$\pm$1.83 62.59$\pm$0.61 62.81$\pm$0.65 3.84 $\pm$ 3.84
$\mathcal{S}_{1}$ 64.10$\pm$0.28 64.10$\pm$0.28 63.74$\pm$0.91 63.57$\pm$0.79 64.10$\pm$0.28 44.93 $\pm$ 29.42
$\mathcal{S}_{2}$ 64.73$\pm$0.79 64.54$\pm$0.63 65.52$\pm$0.20 63.86$\pm$0.32 65.17$\pm$0.20 45.62 $\pm$ 29.86
$\mathcal{S}_{3}$ 66.26$\pm$0.28 66.07$\pm$0.33 66.31$\pm$0.21 65.40$\pm$0.40 66.31$\pm$0.21 19.91 $\pm$ 30.41
$\mathcal{S}_{4}$ 66.89$\pm$1.66 67.33$\pm$0.39 50.26$\pm$27.14 66.85$\pm$0.83 67.57$\pm$0.23 29.42 $\pm$ 31.82
$\mathcal{S}_{5}$ 61.98$\pm$14.03 68.86$\pm$0.18 62.21$\pm$20.31 53.91$\pm$18.89 68.98$\pm$0.18 22.17 $\pm$ 0.00
@@
$\mathcal{S}_{0}$ 54.01$\pm$0.54 54.03$\pm$0.06 53.98$\pm$0.58 50.28$\pm$5.35 54.18$\pm$0.52 2.30 $\pm$ 3.52
$\mathcal{S}_{1}$ 55.05$\pm$0.39 55.08$\pm$0.04 54.33$\pm$1.37 54.21$\pm$0.94 55.81$\pm$0.40 37.99 $\pm$ 24.95
$\mathcal{S}_{2}$ 55.32$\pm$2.13 55.78$\pm$0.06 56.27$\pm$0.24 52.95$\pm$4.04 56.27$\pm$0.24 28.12 $\pm$ 28.12
$\mathcal{S}_{3}$ 57.81$\pm$0.31 57.33$\pm$0.03 57.82$\pm$0.28 57.19$\pm$0.37 57.85$\pm$0.25 23.13 $\pm$ 28.33
$\mathcal{S}_{4}$ 57.64$\pm$4.43 59.05$\pm$0.04 59.54$\pm$0.29 57.98$\pm$3.21 59.54$\pm$0.29 16.03 $\pm$ 24.94
$\mathcal{S}_{5}$ 49.87$\pm$16.51 60.72$\pm$0.09 61.15$\pm$0.19 55.39$\pm$6.39 61.15$\pm$0.19 22.17 $\pm$ 0.00
@@
SNN-2 SCNN-2 SCNN-3 Bi-SCNN-3 Bi-SCNN-2 MPNN-2